# The Calibr-A-Ton: a novel method for calorimeter energy calibration

### v1

@@ 3.5.2 Resolution effects
where the term in $a,b,c$ represents the stochastic, noise and constant terms, respectively, that are added in quadrature. Only the stochastic component is modeled through GEANT4 in our simulation. The effect of the noise term on the electromagnetic shower resolution in HGCAL has been estimated to be negligible [5] and for that reason is not included here either. The shower energy resolution in the presence or the absence of energy biases is shown in Figure 7. The points are fitted with functions of the form of the one of Eq. 10; the extraction of a $c\neq 0$ parameter in the unbiased case shows that the simulation is already affected by a constant term ($c=0.01$), albeit negligible compared to the stochastic term ($a=0.21\;\sqrt{\mathrm{GeV}}$) in the energy range considered. The fact that the constant term is much amplified ($c=0.08$) when an energy bias, of the form described in Subsection 3.5.1, is applied, means that the energy bias acts as an additional source of uncertainty on the shower energy measurement. In this paper, we assess the capability of the method to mitigate the impact of biases, such that the energy resolution after calibration, in the presence of biases, is as close as possible to the uncalibrated resolution, in the absence of biases.
Figure 7: Resolution on the shower energy (estimated by the root mean square RMS of the response histogram, divided by the median generator energy), as a function of the generator energy. Functions of the form $f(E_{\mathrm{gen}})=\sqrt{a^{2}/E_{\mathrm{gen}}+c^{2}}$ are fitted to these points. As expected, the unbiased resolution (in dark green) presents a negligible constant term and is dominated by the stochastic term ($a=0.21\;\sqrt{\mathrm{GeV}}$). As illustrated by the blue points, the introduction of the energy biases outlined in subsection 3.5.1 leads to a substantial constant term ($c=0.08$) but leaves the stochastic term nearly unchanged ($a=0.22\;\sqrt{\mathrm{GeV}}$), as expected.
@@ 4 Performance
The method described in Section 2, the geometry and data sample described in Section 3, are used to extract the calibration constants described in Subsection 3.4 for either methods (per-layer calibration and Calibr-A-Ton) in the presence of the biases introduced in Subsection 3.5.1. In this section, the resulting performance coming from either method are compared. The input dataset is split between a training sample, containing $N_{\mathrm{train}}=3.5\times 10^{5}$ electrons, and a testing sample, containing $N_{\mathrm{test}}=1.5\times 10^{5}$ electrons. The samples are further split into batches of size $N=500$ electrons. The calibration constants are all initialized at one. The JAX open-source library is used to perform an automatic differentiation of a function encoding the loss function of Eq. 3. The value of the loss function as a function of the batch number is shown for either method on Figure 8 (left).
Figure 8: (Left) Value of the loss function of Eq. 3 as a function of the batch number for the per-layer calibration (in magenta) and for Calibr-A-Ton (in blue). The training losses (dashed line) are evaluated as the average of the losses of the previous 14 training batches or 14 weight updates, corresponding in total to 7000 training electrons. The testing losses (solid line) are computed once every 14 batch. (Right) Value of the learning rate vs. the batch number, a process called “annealing”. The same learning rate values are used for either methods.
The distributions of the response, as defined as in Eq. 5, for the inclusive testing sample are shown for the different conditions and for the different training methods in Figure 9. With respect to the raw response of the calorimeter, the application of the biases distorts the response, and in particular generates a double-peak structure, which is not absorbed using the per-layer calibration. On the other hand, the calibration constants derived with the Calibr-A-Ton method allow for an almost full absorption of the bias, providing a response very similar to the situation without bias. This is confirmed when looking at the response in bins of the generator electron energy, as shown in Figure 10.
Figure 9: Histograms of the response, as defined in Eq. 5, for electron showers (the absorber correction is included), when: no bias is introduced (green); the bias is applied but no calibration constants are derived (light blue); the bias is applied and the calibration constants from the per-layer calibration are included (gray); the bias is applied and the calibration constants from Calibr-A-Ton are included (magenta).
@@
The resolution, in bins of the electron generator energy, shown in Figure 11 for the different cases, confirms these conclusions. On the one hand, it can be observed that the per-layer calibration is unable to compensate for the biases introduced, and presents a resolution and a constant term ($c=0.07$) comparable to the uncalibrated, biased case ($c=0.08$). On the other hand, applying the calibration constants derived with the Calibr-A-Ton method leads to a resolution similar to the unbiased case, in the whole energy range. Almost identical stochastic and constant terms are extracted: $a=0.22\;\sqrt{\mathrm{GeV}}$, $c=0.02$ for Calibr-A-Ton when they were $a=0.21\;\sqrt{\mathrm{GeV}}$, $c=0.01$ in the unbiased case. Table 1 summarizes these results. The achieved performance illustrates the potential of the Calibr-A-Ton method. It enables the derivation of calibration constants that depend on all relevant dimensions of the problem – in this application, the 3D coordinates and energies of the cells. Additionally, it employs a custom binning approach that adapts to the available statistics along these dimensions, ensuring optimal data utilization – since each shower typically deposits energy in hundreds of cells, the method maximizes the statistical information available for determining numerous calibration constants. These features make Calibr-A-Ton a promising and highly suitable method for the calibration of high-granularity calorimeters such as the CMS HGCAL.
@@
Table 1: Value of the resolution parameters for the stochastic term ($a$) and the constant term ($c$) in the different cases under study.
@@ 5 Conclusion
This paper describes a novel method for the calibration of calorimeters, Calibr-A-Ton, that bypasses some of the major constraints found in historical calibration procedures. To test the implementation of the method, based on differential programming, and its performance, the simulation of a high granularity calorimeter, with one electromagnetic and two hadronic sections, is used. Electrons of known energies are allowed to shower in this detector, typically leaving energy in hundreds of cells. Known biases are then introduced on the cells energy response, to emulate the resolution effects that typically affect calorimeters. The results clearly show that, contrary to a simple approach based on layer-per-layer calibration, the Calibr-A-Ton method is able to meaningfully compensate for these biases and to nearly recover the original detector resolution. When fitting the resolution curve of the shower energy, almost identical stochastic and constant terms are extracted. On top of the performance achieved, it should be noted that Calibr-A-Ton is also more versatile than other methods, as it has the potential to be used in several contexts: simulation-based calibration (as is the case here), test beams, and in situ calibrations using physics signals or auxiliary measurements by another subdetector, such as the tracker.
This work represents a proof-of-principle of the Calibr-A-Ton method in the successful extraction of calibration constants for the electromagnetic section of a high-granularity calorimeter similar to the CMS HGCAL. Subsequent work will explore the calibration of the hadronic sections using simulated charged pion showers; the introduction of biases directly in the GEANT4 simulation (material density fluctuations, presence of dead material, etc.); the effects of a different choice of binning of the calibration constants on the performance; etc.

### v2

@@ 4.2 Resolution effects
where the terms associated with $a,b,c$ represent the stochastic, noise and constant terms, respectively, that are added in quadrature. Only the stochastic component is modeled through GEANT4 in our simulation. The effect of the noise term on the electromagnetic shower resolution in HGCAL has been estimated to be negligible [8] and for that reason is not included here either ($b=0$). The shower energy resolution in the presence or the absence of energy biases is shown in Figure 7. The points are fitted with functions of the form of the one of Eq. 10. Thanks to the large statistics involved in the fits, the relative uncertainties on the $a,c$ parameters extracted are always less one the per-mil and are therefore not reported. The unbiased case shows that the simulation is already affected by a constant term ($c=0.01$), albeit negligible compared to the stochastic term ($a=0.21\;\sqrt{\mathrm{GeV}}$) in the energy range considered. The fact that the constant term is much amplified ($c=0.08$) when an energy bias, of the form described in subsection 4.1, is applied, means that the energy bias acts as an additional source of uncertainty on the shower energy measurement. In this paper, we assess the capability of the method to mitigate the impact of biases, such that the energy resolution after calibration, in the presence of biases, is as close as possible to the uncalibrated resolution, in the absence of biases.
Figure 7: Resolution on the shower energy, estimated by the root mean square, RMS, of the response histogram divided by the median generator energy, as a function of the generator energy. Functions of the form $f(E_{\mathrm{gen}})=\sqrt{a^{2}/E_{\mathrm{gen}}+c^{2}}$ are fitted to these points. As expected, the unbiased resolution (in dark green) presents a negligible constant term and is dominated by the stochastic term ($a=0.21\;\sqrt{\mathrm{GeV}}$). As illustrated by the blue points, the introduction of the energy biases outlined in subsection 4.1 leads to a substantial constant term ($c=0.08$) but leaves the stochastic term nearly unchanged ($a=0.22\;\sqrt{\mathrm{GeV}}$), as expected.
@@ 5 Performance
The method described in Section 2, and the geometry and data sample described in Section 3, are used to extract the calibration constants described in subsection 3.4 for either methods (per-layer calibration and Calibr-A-Ton) in the presence of the biases introduced in subsection 4.1. In this section, the resulting performance coming from either method are compared. The input dataset is split between a training sample, containing $N_{\mathrm{train}}=3.5\times 10^{5}$ electrons, and a testing sample, containing $N_{\mathrm{test}}=1.5\times 10^{5}$ electrons. The samples are further split into batches of size $N=500$ electrons. The calibration constants are all initialized at 1.0. The JAX open-source library [9] is used to perform an automatic differentiation of a function encoding the loss function of Eq. 3. The value of the loss function as a function of the batch number is shown for either method on Figure 8 (left).
Figure 8: (Left) Value of the loss function of Eq. 3 as a function of the batch number for the per-layer calibration (in magenta) and for Calibr-A-Ton (in blue). The training losses (dashed line) are evaluated as the average of the losses of the previous 14 training batches or 14 weight updates, corresponding in total to 7000 training electrons. The testing losses (solid line) are computed once every 14 batches. (Right) Value of the learning rate vs. the batch number, a process called “annealing”. The same learning rate values are used for either methods.
The distributions of the response, as defined in Eq. 5, for the inclusive testing sample are shown for the different conditions and for the different training methods in Figure 9. With respect to the raw response of the calorimeter, the application of the biases distorts the response, and in particular generates a double-peak structure, which is not absorbed using the per-layer calibration. On the other hand, the calibration constants derived with the Calibr-A-Ton method allow for an almost full absorption of the bias, providing a response very similar to the situation without bias. This is confirmed when looking at the response in bins of the generator electron energy, as shown in Figure 10.
Figure 9: Histograms of the response, as defined in Eq. 5, for electron showers, when: no bias is introduced (green); the bias is applied but no calibration constants are derived (light blue); the bias is applied and the calibration constants from the per-layer calibration are included (gray); the bias is applied and the calibration constants from Calibr-A-Ton are included (magenta).
@@
The resolution, in bins of the electron generator energy, shown in Figure 11 for the different cases, confirms these conclusions. On the one hand, it can be observed that the per-layer calibration is unable to compensate for the biases introduced, and presents a resolution and a constant term ($c=0.07$) comparable to the uncalibrated, biased case ($c=0.08$). On the other hand, applying the calibration constants derived with the Calibr-A-Ton method leads to a resolution similar to the unbiased case, in the whole energy range. Similar stochastic and constant terms are extracted: $a=0.22\;\sqrt{\mathrm{GeV}}$, $c=0.02$ for Calibr-A-Ton when they were $a=0.21\;\sqrt{\mathrm{GeV}}$, $c=0.01$ in the unbiased case. Table 1 summarizes these results. The achieved performance illustrates the potential of the Calibr-A-Ton method. It enables the derivation of calibration constants that depend on all relevant dimensions of the problem – in this application, the 3D coordinates and energies of the cells. Additionally, it employs a custom binning approach that can be adapted to the available statistics along these dimensions. Finally, since each shower typically deposits energy in hundreds of cells, the method maximizes the statistical information available for determining numerous calibration constants. These features make Calibr-A-Ton a promising and highly suitable method for the calibration of high-granularity calorimeters such as the CMS HGCAL.
@@
Table 1: Value of the resolution parameters for the stochastic term ($a$) and the constant term ($c$) in the different cases under study. Uncertainties on those parameters are negligible.
@@ 6 Conclusion
This paper describes a novel method for the calibration of calorimeters, Calibr-A-Ton, that bypasses some of the major constraints found in historical calibration procedures. To test the implementation of the method, based on differential programming, and its performance, the simulation of a high granularity calorimeter, with one electromagnetic and two hadronic sections, is used. Simulated electrons of known energies are generated and produce shower in this detector, typically leaving energy in hundreds of cells. Known biases are then introduced on the cells energy response, to emulate the resolution effects that typically affect calorimeters. The results clearly show that, contrary to a simple approach based on layer-per-layer calibration, the Calibr-A-Ton method is able to meaningfully compensate for these biases and to nearly recover the original detector resolution. When fitting the resolution curve of the shower energy, almost identical stochastic and constant terms are extracted. On top of the performance achieved, it should be noted that Calibr-A-Ton is also more versatile than other methods, as it has the potential to be used in several contexts: simulation-based calibration (as is the case here), test beams, and in situ calibrations using physics signals or auxiliary measurements by another subdetector, such as the tracker.
This work represents a proof-of-principle of the Calibr-A-Ton method in the successful extraction of calibration constants for the electromagnetic section of a high-granularity calorimeter similar to the CMS HGCAL. Subsequent work will explore the calibration of the hadronic sections using simulated charged pion showers, the introduction of biases directly in the GEANT4 simulation (material density fluctuations, presence of dead material, etc.), and the effects of a different choice of binning of the calibration constants on the performance.